OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

i will give you a medal

OpenStudy (mathstudent55):

Let x = number of one-step equations Let y = number of two-step equations "The number of one-step equations and two-step equations who have been eliminated today is equal to 1070." This means: x + y = 1070 "three times the number of one-step equations minus twice the number of two-step equations is equal to 110" means: 3x - 2y = 110 You have a system of equations: x + y = 1070 3x - 2y = 110 Multiply the first equation by 2, and add to the second equation: 2x + 2y = 2140 3x - 2y = 110 ---------------(add) 5x = 2250 x = 450 Answer: 450 one-step equations